p53. A basketball has a mass of 570g. Moving to the right and heading downward at an angle of 30o to the vertical, it hits the floor with a speed of 5m/s and bounces up with nearly the same speed, again moving to the right at an angle of 30o to the vertical. What was the momentum change (Delta P)?
Answer #1

Bons AP Is u= 5 in 301 - 500s3oj V = 5 Sin 30i + 5 cos 30j AP = mv-onu > m(V-u) ~0.57023 * (5 Sn301+5 cos3oj. - 55n30 it scos